# Recurrent Diffusion for Large-Scale Parameter Generation

### v1

@@ 4.2 Results of Large-Scale Parameter Generation
In this section, we present the results of our approach across a range of tasks including classification, semantic segmentation, object detection&instance segmentation, and language tasks. As most previous works encounter the out-of-memory issue at million-scale parameter generation, we mainly compare with the results from the trained networks, which we denote as ‘original’.
Results on ImageNet-1K.
Tab. 1 presents performance comparisons across seven architectures on ImageNet-1K [12]. These architectures encompass the ResNet [22], ViT [14], and ConvNeXt [41] series, with parameter counts ranging from 3 to 197 million. Based on the results in Tab. 1, several crucial observations can be made as follows: i) Our approach successfully generates model parameters at hundred-million scales, overcoming the out-of-memory issues faced by previous works [51, 69, 66, 29]. ii) The performances of the generated models are comparable with the original ones.
@@
Table 2: Accuracy comparison of original and generated parameters on ADE20K (176.5M parameters) and COCO (110.9M parameters). All models are built based on ViT-Base [14].
Results on ADE20K and COCO.
We also investigate the generalization of our approach to semantic segmentation as well as object detection and instance segmentation tasks. We choose ADE20K [77] and COCO [39] as our benchmark datasets. For semantic segmentation, following  Zhao et al. [76], we adopt UperNet [72] as the segmentation model and train it on ADE20K to prepare checkpoints. For object detection and instance segmentation, we finetune ViTDet [36] on COCO to collect checkpoints and report the results of mAP Bbox and mAP Seg, respectively. All experiments here are conducted based on ViT-B [14]. Tab. 2 presents the strong generalization of our approach to these two tasks. Specifically, compared to the original models, we achieve comparable or even slightly better results over all the above metrics.
@@
original RPG original RPG original RPG original RPG original RPG original RPG original RPG
@@
Results on commonsense reasoning.
We conduct experiments on language tasks to evaluate the generalization of our approach. We employ DoRA [40], an upgraded version of LoRA [28], to fine-tune LLaMA-7B [67] for commonsense reasoning tasks and save the checkpoints as the training data. We report the results across 7 sub-tasks with rank = 4 and 64 in Tab. 3. The generated models consistently yield results comparable to those of the original ones.
@@ 4.3 Ablation Studies and Analysis
In this section, we examine the influences of key factors on our method. We present the results of the generated ViT-Tiny [14] on the ImageNet-1K [12], unless stated otherwise.
@@
In ViT [14], the position embeddings are learnable by default. Here, we mainly conduct the experiments with three different position embedding manners and show the details as follows:
@@
As shown in Tab. LABEL:tab:pe_ablation, the learnable embeddings perform slightly better than the other two manners. However, we still recommend using fixed position embeddings, as they offer comparable performance while significantly reducing storage requirements compared to the learnable position embedding scheme.
@@
Considering the differences among various layers, we divide the parameters into tokens within each layer. P-diff [69] directly flattens the parameters into 1-dimensional vectors, while SANE [61] divides the parameters by channel within each layer. We conduct experiments to analyze the 3 strategies mentioned above and compare their results in Tab. LABEL:fig:manner_of_tokenization. Our default strategy achieves better results than the others. Directly flattening results in a single token containing parameters from different layers, which poses challenges for optimization. Tokenizing by the channel may result in excessive padding values for each token, as the number of channels is usually much smaller than the default token size.
@@
We mainly explore three structures of the recurrent model, including LSTM [26], Transformer [68], and Mamba [20]. In Tab. 5, we report the performances, training time, and memory cost of generating ViT-Tiny parameters on ImageNet-1K. All results are obtained on a NVIDIA H100 80G GPU. All three structures can achieve good results. However, considering the training time and memory cost, our default Mamba is the best structure of the recurrent model.
@@
LSTM [26] 75.5 75.2 74.4 16.1 38.1
Transformer [68] 75.0 74.8 74.6 4.2 29.1
Mamba [19] 75.4 75.3 75.2 4.1 27.8
@@
In Tab. 6, we show the results of generated ViT series with tokens of different sizes ranging from 1024 to 16384. The performance of generated models become better as the token size increases. When token is of small size, it will contain limited information that is hard to learn. Despite carrying large volume of information, larger token will lead to substantial memory costs (see in Appendix Fig. 9(a)).
@@
The effect of permutation state.
collected runs org. w/o permu. state w/ permu. state
1 88.2 88.0 88.1
3 88.3 fail 88.1
10 88.5 fail 88.2
@@
Rapid synthesis of large-scale parameters is crucial for evaluating the practicality of our approach. As illustrated in Tab. 8, we present the time cost for generating models of ViT-Base and ConvNeXt-L across various DDIM [64] sampling steps. All results are obtained with a single NVIDIA H100 80G GPU. Our approach shows the capability to generate models within minutes. Notably, even for ConvNeXt-L (197.7 M parameters), we can synthesize the entire parameter within 1.3 minutes. Even with only 20 sampling steps, we can achieve promising results. Meanwhile, the inference memory requirement is approximately 20GB, so RPG can be deployed on NVIDIA RTX 3090 or similar-level GPUs.
@@
In this section, we demonstrate that our method offers significant advantages over simply adding noise to the original models. Following p-diff [69], we choose Intersection of Union (IoU) as the metric for measuring similarity. It compares the agreement of output results from classification models across a large number of samples to evaluate the similarity. We calculate IoU of models from target groups with all the original models and select the maximum IoU value (nearest neighbor) as the measure of similarity.
In Fig. 4, we compare original models adding various levels of noise with models generated by our method in terms of accuracy and similarity. As the noise level increases, the similarity and accuracy of the models both decreases. The points representing our generated models are distributed in the upper left region relative to the area with added noise, indicating that our models can enhance diversity while maintaining accuracy.
Figure 4: The figures show the trade-off between accuracy and similarity. The shaded area includes the approximate range of noise-added checkpoints. These two plots demonstrate the strong accuracy and similarity trade-off of our approach.
@@ 4.4 Comparisons with Previous Methods
We compare our approach with four previous works, i.e., $S_{\text{KDE30}}$ [58], p-diff [69], D2NWG [66], and SANE [61]. As shown in Tab. 9, our approach consistently achieves the best results on various architectures, while previous works are hard to achieve comparable performances as original models. Another key issue is that the previous works usually fail to generate large-scale neural network parameters. We also provide comparisons between RPG and previous works in Tab. 10. RPG demonstrates superior capabilities in all three aspects: scalability, performance, and architecture support.
[…]
params. (M) 0.003 0.011 4.2 11.7 86.6
original 49.0 62.1 95.4 93.9 98.7
$S_{\text{KDE30}}$ [58] 26.9 - OOM OOM OOM
p-diff [69] 48.8 61.9 OOM OOM OOM
D2NWG [66] 38.2 - 82.2 - -
SANE [61] - 57.9 - 68.6 -
RPG 49.0 62.0 89.3 93.6 98.9
Table 9: Our approach obtains the best results across all architectures and largely outperforms previous works. Underline and OOM denote the reproduced results and out-of-memory issue, respectively. The detailed structures of CNN (s) and CNN (m) can be found in Model Zoos [60].
method scalability performance support popular arch.
$S_{\text{KDE30}}$ [58]
p-diff [69]
D2NWG [66]
SANE [61]
RPG
Table 10: Comparison of RPG and existing methods on key aspects: scalability to large models, effectiveness of generated parameters, and support for popular architectures. RPG excels in all three dimensions.
Figure 5: An illustration of our binary embedding strategy and dataset construction. Left: binary embeddings (1022 in total) encode different CIFAR-10 classification tasks, where 1s indicate classes to be classified together (e.g., ’ship’ and ’truck’ in the first example). Right: the dataset consists of parameter-encoding pairs, formed by network parameters with their corresponding binary embeddings. These pairs are split into non-overlapping training and validation sets.
unseen tasks (w.r.t. embeddings) original RPG
@@
5 RPG’s Potential in Unseen Tasks
@@
Table 12: Result comparisons of binary embedding change. The accuracy in the table is for each individual class. RPG can be aware of such change accurately. More results can be found in Appendix B.4.
5.1 Experiment Designs.
Build seen and unseen tasks.
To assess RPG’s capability in generating models for unseen tasks, we construct various binary classification tasks on CIFAR-10. As illustrated in Fig.  5, we encode each task as a 10-bit binary embedding, where each bit corresponds to a CIFAR-10 category. A value of 1 indicates that the corresponding category should be classified as positive (class 1), while 0 indicates negative (class 0). For example, in the third embedding shown, ‘ship’ and ‘truck’ are assigned to class 1, while all other categories belong to class 0.
(a) Original and generated models with identical seen binary embeddings are compared. The three original models exhibit homogeneity, while the generated models display diversity.
@@
Collection of the checkpoints.
Training of RPG.
We only use the 1002 checkpoints from seen tasks to train RPG. Meanwhile, these embeddings are also fed into RPG as conditional inputs of the recurrent model. During the training stage, the checkpoints trained by the unseen binary embeddings are not accessible.
Evaluation details.
We input the unseen binary embeddings to the trained RPG to generate the parameters for unseen tasks. The results of the original and generated unseen models are reported for comparison.
5.2 Results for Unseen Tasks
Performance comparisons.
We compare the results of our approach and original models on unseen binary embeddings in Tab. 11. Considering the space limitation, we randomly select 10 unseen binary embeddings for comparison. Notably, RPG yields commendable performance in these unseen tasks, even without being trained on the specific unseen embeddings. That demonstrates the strong practicality and potential of our approach in generating models under unseen tasks. The results of the remaining unseen binary embeddings and more analysis are shown in Appendix B.4.
Perception of embedding changes.
Visualizations of original and generated model parameters.
@@ 5.2 Results for Unseen Tasks
epoch from scratch finetune RPG initialization + finetune
0 50.0 52.9 94.4
1 61.4 90.3 96.8
5 69.1 96.3 97.3
10 74.9 97.1 97.5
50 86.7 97.7 97.8
Efficiency comparison.
@@ 6 Discussion and Conclusion
Our approach demonstrates promising results in large-scale parameter generation across various vision and language tasks. However, we acknowledge that achieving true ‘AI creating AI’ remains a distant goal. Firstly, while our method shows potential in generating models for unseen tasks, it currently faces limitations in generating parameters for novel model architectures. Secondly, our approach is constrained by modeling parameter relationships within a single task, potentially limiting its practical applicability. More importantly, future work should focus on simultaneously modeling parameter relationships across diverse architectures and tasks. Such an approach could yield a more powerful and versatile parameter generator, potentially advancing us closer to the ‘AI creating AI’ era. We hope our approach will inspire future research in this field.
@@ Acknowledgments.
We sincerely thank Zhiyuan Liang, Gongfan Fang, Xuanlei Zhao, Zangwei Zheng, Ziheng Qin, Tianlong Chen, and Zhangyang Wang for valuable discussions and feedbacks. This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-08-008).
@@ Appendix A Discussion with More Related Works
A.1 Discussion with HyperRepresentations.
We mainly compare with three HyperRepresentation methods [58, 61, 59]. These methods use an autoencoder to learn the latent features of trained models, so they call the latent feature HyperRepresentation. This HyperRepresentation is then used for analyzing the model’s performance or characteristics, or for sampling to generate new models or pre-trained parameters.
  • [58] utilizes kernel density estimation (KDE) to sample model parameters on the learned HyperRepresentation space. They also emphasize the importance of layer-wise loss normalization in the learning process of HyperRepresentation. This work achieves parameter generation in small CNNs from Model Zoos [60] with 2864 parameters.
  • [59] focuses on using HyperRepresentation to sample the pre-trained model parameters. They also evaluate the ability of transfer learning by using a trained parameter autoencoder to initialize on unseen dataset. This work can be regarded as a cheap parameter initialization method.
  • [61] utilizes a sequential autoencoder for neural embeddings (SANE) to divide the neural network weights into subsets. Then, an autoencoder processes these subsets in a sliding window. This work can generate the entire parameter of ResNet-18. However, the performance of generated ResNet-18 is poor and exists a large gap with the original trained ResNet-18. For example, the performance of generated ResNet-18 is 68.6%, while the original model can achieve an accuracy of over 90% on CIFAR-10 (see in Tab. 9).

### v2

@@ 3.2 Results of Large-Scale Parameter Generation
On ImageNet-1K.
Tab. 1 presents performance comparisons across seven architectures on ImageNet-1K. The parameter number of these architectures ranges from 3 to 197 million. Several observations can be made as follows: i) Our approach successfully generates model parameters at hundred-million scales, overcoming the out-of-memory issues faced by previous works (Peebles et al., 2022; Wang et al., 2024). ii) The performances of the generated models are comparable with the original ones.
@@
Table 2: Accuracy comparison of original and generated parameters on ADE20K (176.5M parameters) and COCO (110.9M parameters). All models are built based on ViT-Base.
@@
org RPG org RPG org RPG org RPG org RPG org RPG org RPG
@@
On ADE20K and COCO.
For semantic segmentation, following  Zhao et al. (2024), we adopt UperNet (Xiao et al., 2018) as the segmentation model and train it on ADE20K (Zhou et al., 2017) to prepare checkpoints. For object detection and instance segmentation, we finetune ViTDet (Li et al., 2022) on COCO (Lin et al., 2014) to collect checkpoints and report the results of mAP Bbox and mAP Seg, respectively. All experiments here are conducted based on ViT-Base (Dosovitskiy et al., 2021). Tab. 2 presents the strong generalization of RPG to these two tasks. Specifically, compared to the original models, we achieve comparable or even slightly better results over all the above metrics.
@@
On commonsense reasoning.
We employ DoRA (Liu et al., 2024), an upgraded version of LoRA (Hu et al., 2022), to fine-tune LLaMA-7B (Touvron et al., 2023) for commonsense reasoning tasks and save the checkpoints as the training data. We report the results across 7 sub-tasks with rank = 4 and 64 in Tab. 3. The generated models consistently yield results comparable to those of the original ones.
@@ 3.3 Ablation Studies and Analysis
In this section, we present the results of the generated ViT-Tiny on the ImageNet-1K, unless stated otherwise.
@@
In ViT, the position embeddings are learnable by default. Here, we mainly conduct the experiments with three different position embedding manners and show the details as follows:
@@
As shown in Tab. LABEL:tab:pe_ablation, the learnable embeddings perform slightly better than the other two manners. However, we still recommend using fixed position embeddings, as they offer comparable performance while significantly reducing storage requirements compared to the learnable one.
@@
Considering the differences among various layers, we divide the parameters into tokens within each layer. P-diff (Wang et al., 2024) flattens the parameters into 1D vectors, while SANE (Schürholt et al., 2024) divides the parameters by channel within each layer. We compare the results of these 3 strategies in Tab. LABEL:fig:manner_of_tokenization. Our default strategy achieves better results than the others. Flattening results in a single token containing parameters from different layers, which poses challenges for optimization. Tokenizing by the channel may result in excessive padding values for each token, as the channel number is usually much smaller than the token size.
@@
We mainly explore three structures of the recurrent model, including LSTM (Hochreiter et al., 1997), Transformer (Vaswani, 2017), and Mamba (Gu & Dao, 2024). In Tab. 5, we report the performances, training time, and memory cost of generating ViT-Tiny parameters on ImageNet-1K. All three structures can achieve good results. However, considering the training time and memory cost, our default Mamba is the best structure of the recurrent model.
@@
LSTM 75.5 75.2 74.4 16.1 38.1
Transformer 75.0 74.8 74.6 4.2 29.1
Mamba 75.4 75.3 75.2 4.1 27.8
@@
In Tab. 6, we show the results of generated ViT series with tokens of different sizes ranging from 1024 to 16384. The performance of generated models become better as the token size increases. When token is of small size, itcontains limited information that is hard to learn. Despite carrying large volume of information, larger token leads to substantial memory costs (see in Appendix Fig. 9(a)).
@@
Rapid synthesis of large-scale parameters is crucial for evaluating the practicality of our approach. As illustrated in Tab. 7, we present the time cost for generating models of ViT-Base and ConvNeXt-L across various DDIM (Song et al., 2020) sampling steps. All results are obtained with a single NVIDIA H100 80G GPU. Our approach shows the capability to generate models within minutes. Notably, even for ConvNeXt-L (197.7 M parameters), we can synthesize the entire parameter within 1.3 minutes. Even with only 20 sampling steps, we can achieve promising results. Meanwhile, the inference memory requirement is approximately 20GB, so RPG can be deployed on NVIDIA RTX 3090 or similar-level GPUs.
@@
We compare RPG with adding noise and model interpolation. Following p-diff (Wang et al., 2024), we use IoU to measure model similarity by comparing outputs across samples, selecting the maximum IoU with original models as our metric.
In Fig. 4, as the noise level increases, both the similarity and accuracy of the models decrease. The points representing our generated models are distributed in the upper left region relative to the other conditions, indicating that our models can enhance diversity while maintaining accuracy. Furthermore, our method demonstrates greater diversity in comparison to trivial interpolation methods.
Figure 4: The figure shows the trade-off between accuracy and similarity with ViT-Tiny on ImageNet-1K. The shaded area includes the approximate range of noise-added checkpoints. This plot demonstrates the strong trade-off between accuracy and similarity and highlights our advantages over trivial interpolation.
@@ 3.4 Comparisons with Previous Methods
We compare RPG with four previous works, i.e., $S_{\text{KDE30}}$ (Schürholt et al., 2022a), p-diff (Wang et al., 2024), D2NWG (Soro et al., 2024), and SANE (Schürholt et al., 2024). We report the original and generated performances for comparison. As shown in Tab. 8, among all compared methods, only our approach consistently achieves the highest results and comparable results to original models across various architectures. Another key issue is that the previous works usually fail to generate large-scale neural network parameters. (More discussion is in Appendix A.)
method CNN (s) CNN (m) ResNet-18 ViT-Base
params. (M) 0.003 0.011 11.7 86.6
$S_{\text{KDE30}}$ 26.9 46.1 - OOM OOM
p-diff 48.8 49.0 61.9 62.1 OOM OOM
SANE - 57.9 57.2 68.6 85.5 -
D2NWG 38.2 44.7 58.8 57.2 94.6 94.6 -
RPG 49.0 49.0 62.0 62.1 95.1 95.3 98.9 98.7
Table 8: Our approach obtains the best results across all architectures and largely outperforms most previous works. The subscripts denote the average accuracy of corresponding original models. Underline denotes the reproduced results. And OOM denotes out-of-memory issue. The detailed structures of CNN (s) and CNN (m) can be found in Model Zoos (Schürholt et al., 2022).
Figure 5: An illustration of our binary embedding strategy and dataset construction. Left: binary embeddings (1022 in total) encode different CIFAR-10 classification tasks, where 1s indicate classes to be classified together (e.g., ‘ship’ and ‘truck’ in the first example). Right: the dataset consists of parameter-encoding pairs, formed by network parameters with their corresponding binary embeddings. These pairs are split into non-overlapping training and validation sets.
4 RPG Generalizes to Unseen Tasks
4.1 Experiment Designs.
Build seen and unseen tasks.
To assess RPG’s capability for unseen tasks, we construct various binary classification tasks on CIFAR-10. As shown in Fig.  5, we encode each task as a 10-bit binary embedding, where each bit corresponds to a CIFAR-10 category. 1 indicates that the corresponding category should be classified as positive (class 1), while 0 indicates negative (class 0). For example, in the third embedding shown, ‘ship’ and ‘truck’ are assigned to class 1, while all other categories belong to class 0.
Collection of the checkpoints.
Training and evaluation of RPG.
We only use the 1002 checkpoints from seen tasks to train RPG. Meanwhile, these embeddings are also fed into RPG as conditional inputs of the recurrent model. During the training stage, the checkpoints trained by the unseen binary embeddings are not accessible. When evaluating, we input the unseen binary embeddings to the trained RPG to generate the parameters for unseen tasks. The results of the original and generated unseen models are reported for comparison.
4.2 Results for Unseen Tasks
unseen tasks (embeddings) original RPG
@@ 4.2 Results for Unseen Tasks
Performance comparisons.
We compare the results of our approach and original models on unseen binary embeddings in Tab. 9. Considering the space limitation, we randomly select 10 unseen binary embeddings for comparison. Notably, RPG yields commendable performance in these unseen tasks, even without being trained on the specific unseen embeddings. That demonstrates the strong practicality and potential of our approach in generating models under unseen tasks. The results of the remaining unseen binary embeddings and more analysis are shown in Appendix C.2.
@@
Table 10: Comparisons of binary embedding changes. The accuracy is for each single class. RPG can be aware of such changes accurately.
Perception of embedding changes.
(a) Original and generated models with identical seen binary embeddings are compared. The three original models exhibit homogeneity, while the generated models display diversity and maintain high accuracy across all three models.
@@
Visualizations of model parameters.
@@
5 Related Works
In this section, we mainly introduce recurrent models and previous works of parameter generation.
Recurrent models.
Recurrent neural networks (RNNs) were first proposed to process sequential data. To tackle the vanishing gradient problem in early RNNs, long short-term memory (LSTM) (Hochreiter, 1991, 1997) was introduced. Recently, transformer-based models (Vaswani, 2017) exhibits excellent potential in sequential data processing, due to their parallelized training and scalability, such as linear attentions (Wang et al., 2020; Choromanski et al., 2020), RWKV (Peng et al., 2023), Mamba (Gu & Dao, 2023; Dao & Gu, 2024), and xLSTM (Beck et al., 2024).
Parameter generation.
The core idea of parameter generation is to learn the distribution of trained parameters. Stochastic neural networks (Sompolinsky et al., 1988; Bottou et al., 1991; Graves, 2011) and Bayesian neural networks (Neal, 2012; Kingma & Welling, 2013; Gal & Ghahramani, 2016) model the priors or probability distributions over the parameters. However, these methods are limited by their generality to large-scale parameter generation. HyperNetworks (Ha et al., 2017), i.e. is proposed to generate various architectures’ parameters for a larger network. Smash (Brock et al., 2018) extends the range of architectures via a memory read-writes scheme. With the development of diffusion models, many works (Peebles et al., 2022; Chou et al., 2023; Erkoç et al., 2023; Wang et al., 2024; Soro et al., 2024; Lin et al., 2024; Li et al., 2024b; Jin et al., 2024; Peebles et al., 2022) adopt diffusion models to generate neural network parameters. Hyper-Representations  (Schürholt et al., 2022a, b, 2024) use an autoencoder to capture the latent distribution of trained models. COND P-DIFF (Jin et al., 2024) and Tina (Li et al., 2024b) introduce text-controlled parameter generation method. Unfortunately, the above methods have a common drawback: can not generate large-scale parameters, such as whole parameters of ResNet, ViT, ConvNeXt, or LoRA. Therefore, our approach brings new inspiration to the field of parameter generation.
@@ 6 Discussion and Conclusion
Our approach demonstrates promising results in large-scale parameter generation across various vision, language and unseen tasks. However, we acknowledge that achieving true ‘AI generating AI’ remains a distant goal. Firstly, while our method shows potential in generating models for unseen tasks, it currently faces limitations in generating parameters for novel model architectures. Secondly, our approach is constrained by modeling parameter relationships within a single task, potentially limiting its practical applicability. More importantly, future work should focus on simultaneously modeling parameter relationships across diverse architectures and tasks. Such an approach could yield a more powerful and versatile parameter generator, potentially advancing us closer to the ‘AI generating AI’ era. We hope our approach will inspire and encourage future research in neural network parameter generation.
@@ Acknowledgments.
We sincerely thank Zhiyuan Liang, Zhuang Liu, Gongfan Fang, Xuanlei Zhao, Zangwei Zheng, Ziheng Qin, and Tianlong Chen for valuable discussions and feedbacks. This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-08-008).
@@ Appendix A Discussion with More Related Works
A.1 Overview of comparisons
We compare RPG and four other methods from three aspects: scalability, performance, and generalization. Only our RPG excels in all three aspects simultaneously.
method scalability performance generalization
$S_{\text{KDE30}}$
p-diff
SANE
D2NWG
RPG
Table 11: Comparison of RPG and existing methods on key aspects: scalability, performance, and generalization.
A.2 Discussion with Hyper-Representations
We mainly compare with three HyperRepresentation methods (Schürholt et al., 2022a, 2024, b). These methods use an autoencoder to learn the latent features of trained models, so they call the latent feature HyperRepresentation. This HyperRepresentation is then used for analyzing the model’s performance or characteristics, or for sampling to generate new models or pre-trained parameters.
  • (Schürholt et al., 2022a) utilizes kernel density estimation (KDE) to sample model parameters on the learned HyperRepresentation space. They also emphasize the importance of layer-wise loss normalization in the learning process of HyperRepresentation. This work achieves parameter generation in small CNNs from Model Zoos (Schürholt et al., 2022) with 2864 parameters.
  • (Schürholt et al., 2022b) focuses on using HyperRepresentation to sample the pre-trained model parameters. They also evaluate the ability of transfer learning by using a trained parameter autoencoder to initialize on unseen dataset. This work can be regarded as a cheap parameter initialization method.
  • (Schürholt et al., 2024) divides the neural network weights into subsets and utilizes a sequential autoencoder for neural embeddings (SANE) on there subsets with a sliding window. This work can generate the entire parameter of ResNet-18. However, its generation process does not directly derive parameters from noise. Instead, it relies on a half-trained model for Kernel Density Estimation (KDE) sampling.